# Geodesic structure of spacetime near singularities

### v1

@@ IV Discussion
In this work, we have derived a series representation for Synge’s world function for FLRW and Bianchi type I spacetime. Interestingly, the key structural properties of the world function are captured through integrals of $1/a(t)^{2n}$, $n\in\mathbb{N}$, between the end-points of a given timelike geodesic. Such dependence was also observed by Buchdahl long back Buchdahl and Warner (1980) in expansion of $\sqrt{\Omega}$, but this expansion turns out to be bad near singularities due to spurious, unphysical divergences. Our representation, on the other hand, gives well-defined limits if one of the points is singular.
As an important concomitant, we have computed the singular and coincidence limits of the van Vleck determinant for FLRW (matter-dominated and radiation) and Bianchi class of spacetimes, focussing in particular on the Kasner limit of the Schwarzshild near singularity. The van Vleck determinant shows a divergent power law behaviour in the singular limit for matter-dominated FLRW and Kasner cases, whereas a logarithmic divergence for the radiation-dominated FLRW case. We provide some further insights into the geodesic structure close to singularities by investigating the behavior of extrinsic curvature for the equi-geodesic ($\Omega=$ constant) surfaces anchored at a point close to the singularity. The trace of extrinsic curvature gives us the expansion of $\Omega=$ constant surfaces, and we compare this with the expansion of t=t=constant surface; the differences are interesting and shown in Fig. 2. The contrasting behaviour of coincidence and singularity limits of biscalars suggests that the spacetime neighbourhood near a singularity is qualitatively different from a regular region of spacetime. This can have implications for small scale structure of quantum spacetime. Several recent works have demonstrated the important role of bi-tensors in characterizing small scale structure of spacetime; Kothawala (2013),Kothawala (2023a),Kothawala (2023b). It is a major consensus that the Planck length $\ell_{p}$ acts as a universal regulator at small scales and gives rise to non-trivial relics Kothawala (2023b).
@@ Appendix A Calculational details in FLRW spacetime
we can reduce the solution of calculating proper time to quadrature Roberts (1993) and get the following expressions,
@@
where $F$ and $G$ are defined in eqns (56) and (51). Let us now consider $j=1$ term explicitly, The expansion for $G$ is,

### v2

@@ IV Discussion
In this work, we have derived a series representation for Synge’s world function for FLRW and Bianchi type I spacetime. Interestingly, the key structural properties of the world function are captured through integrals of $1/a(t)^{2n}$, $n\in\mathbb{N}$, between the end-points of a given timelike geodesic. Such dependence was also observed by Buchdahl long back [Buchdahl (1972) in expansion of $\sqrt{\Omega}$, but this expansion turns out to be bad near singularities due to spurious, unphysical divergences. Our representation, on the other hand, gives well-defined limits if one of the points is singular.
As an important concomitant, we have computed the singular and coincidence limits of the van Vleck determinant for FLRW (matter-dominated and radiation) and Bianchi class of spacetimes, focussing in particular on the Kasner limit of the Schwarzshild near singularity. The van Vleck determinant shows a divergent power law behaviour in the singular limit for matter-dominated FLRW and Kasner cases, whereas a logarithmic divergence for the radiation-dominated FLRW case. We provide some further insights into the geodesic structure close to singularities by investigating the behavior of extrinsic curvature for the equi-geodesic ($\Omega=$ constant) surfaces anchored at a point close to the singularity. The trace of extrinsic curvature gives us the expansion of $\Omega=$ constant surfaces, and we compare this with the expansion of t=t=constant surface; the differences are interesting and shown in Fig. 2. The contrasting behaviour of coincidence and singularity limits of biscalars suggests that the spacetime neighbourhood near a singularity is qualitatively different from a regular region of spacetime. This can have implications for small scale structure of quantum spacetime. Several recent works have demonstrated the important role of bi-tensors in characterizing small scale structure of spacetime; Kothawala (2013),Kothawala (2023a),Kothawala (2023b), and our work provides important tools to study this structure near singularity, thereby providing a window into quantum effects near singularity.
@@ Appendix A Calculational details in FLRW spacetime
we can reduce the solution of calculating proper time to quadrature (see Roberts (1993)) and get the following expressions,
@@
where $F$ and $G$ are defined in Eqs. (59) and (54). Let us now consider $j=1$ term explicitly, The expansion for $G$ is,